# Test-time Correction with Human Feedback: An Online 3D Detection System via Visual Prompting

### v2

@@ 4.3 Robustness of Visual Prompts
Robustness over Arbitrary Visual Prompt Views. We also investigate the capability of TTC system in handling visual prompts of target objects across scenes and times. In Figure 7, we study whether our system can successfully associate objects with its visual prompts from arbitrary frames. Specifically, for each video clip of the nuScenes validation set, we use the image patch of target objects in the first frame as visual prompts, then detect and track target objects in subsequent frames, and compute the recall rate for evaluation with arbitrary views. Figure 7 presents the results of TTC-MonoDETR, showing that despite significant differences in viewpoints and object poses between frames, the recall rate does not drop dramatically as the ego vehicle moves. This highlights the robustness of the TTC detectors in handling visual prompts with arbitrary views across scenes and times, indicating that the TTC system can effectively process human feedback that may involve temporal delays.
4.4 More Ablation Studies
Due to limited space, we refer readers to Appendix D of the appendix for extensive ablation studies. These studies include: (1.) the ablations on core components in visual prompt alignment in Section D.1 and Section D.2, (2.) the evaluation of visual prompt alignment’s impact on ensuring instance awareness in Section Section D.3, (3.) the examination of statistics regarding the visual prompt buffer during online inference in Section D.4, (4.) the assessment of visual prompt robustness against user perturbations and reduced user feedback collection frequency in Section D.5 and Section D.6, (5.) the comparison highlighting the superiority of our OA module over the combination of 2D single-object trackers with 3D heads for generating 3D tracking in Section Section D.7, and more.
@@ 5 Conclusion
For model design, while allowing user interaction provides reliability, it introduces potential risks. Requiring frequent human feedback may be a challenge in current L3 autonomous driving paradigms. Though reducing the frequency of feedback can alleviate this pressure without compromising performance (See appendix Section D.6), there is room for improvement. Future work could integrate robust tracking methods or leverage advanced hardware, such as eye-tracking, to simplify the feedback collection.
@@ Acknowledgement
OpenDriveLab is the autonomous driving team affiliated with Shanghai AI Lab. This work was supported by National Key R&D Program of China (2022ZD0160104), NSFC (62206172), and Shanghai Committee of Science and Technology (23YF1462000). We thank team members from OpenDriveLab for valuable feedback along the project. Special thanks to Chengen Xie for providing data support, and Jiahui Fu for sharing insights on MV2D.
@@ Appendix A Discussions
We emphasize, in this paper, we do not propose the online system to replace the traditional offline data loop. As illustrated in Fig. 2 of the main paper, these two systems address different aspects of autonomous driving. The offline system remains crucial for enhancing the capabilities of the base perception model through development; while our online TTC system further enables the deployed frozen model in vehicles to promptly rectify dangerous driving behaviors caused by unrecognized objects on the road. With improved offline-trained detectors, the TTC can effectively correct more online errors during test-time inference without re-training, as detailed in Table 1 of the main paper.
@@
Compared to text prompts, visual prompts provide a more natural and accurate description of the target object. In contrast, verbal descriptions can be ambiguous to convey instance-level features, leading to an inaccurate understanding of the missing objects. Second, text promptable models are often combined with LLMs with high latency and are thus unavailable for autonomous driving deployment, as discussed in Section D.8.
Box and point prompts are less convenient than visual prompts when dealing with stream data. If missing occurs, these single-frame prompts require users to provide feedback at every frame, which is unfeasible in real-world applications. Compared to box and point prompts, visual prompts are robust across different scenes and timestamps, one single-frame visual prompt is enough for detecting and tracking in later frames. Furthermore, visual prompts enable 3D detection with pre-defined visual descriptions of target objects, regardless of the sources, styles, poses, etc, as discussed in Section 4.3.
The introduction of novel visual prompts enables real-time, accurate, and continuous error correction of streaming inputs with ”one-click” feedback.
@@
Q4: Why choose 3D detection as the experimental scenarios of TTC? Could the proposed framework be extended to 2D detection or other vision tasks?
TTC represents a general idea to equip deployed systems with the capability of online error rectification, making them more versatile, adaptive, and reliable. This idea can be readily extended to other vision tasks, such as 2D detection, for rapid adaptation of pre-trained models to novel scenarios.
The choice of 3D detection for autonomous driving as the experimental scenario is motivated by the paramount importance of safety for deployed self-driving systems. Without an online correction method, mistakes made by the offline model pose significant safety risks for on-road autonomy.
Therefore, we select the autonomous driving domain as the testbed for the TTC framework, given the critical need for a robust, adaptive online error correction system to ensure the reliability of these safety-critical applications. We mark the extension of TTC to other vision tasks as future works.
Q5: What are potential applications and future directions of TTC?

### v3

@@ 4.3 Robustness of Visual Prompts
Robustness over Arbitrary Visual Prompt Views. We also investigate the capability of TTC system in handling visual prompts of target objects across scenes and times. In Figure˜7, we study whether our system can successfully associate objects with their visual prompts from arbitrary frames. Specifically, for each video clip of the nuScenes validation set, we use the image patch of target objects in the first frame as visual prompts, then detect and track target objects in subsequent frames, and compute the recall rate for evaluation with arbitrary views. Figure˜7 presents the results of TTC-MonoDETR, showing that despite significant differences in viewpoints and object poses between frames, the recall rate does not drop dramatically as the ego vehicle moves. This highlights the robustness of the TTC detectors in handling visual prompts with arbitrary views across scenes and times, indicating that the TTC system can effectively process feedback that may involve temporal delays.
Robustness over Positional Perturbations. Since visual prompts during online inference are derived from user clicks or 2D detectors, they may deviate from the precise 2D center of the target object. To analyze the impact of such positional shifts, we introduce controlled perturbations during test-time inference. Specifically, we apply translation offsets to the ground-truth 2D centers, varying the maximum displacement ratio from 0% to 40%, and evaluate the system’s robustness under these conditions. As shown in Figure˜8, despite increasing perturbations, TTC-MonoDETR exhibits only minor performance degradation, demonstrating strong resilience to positional variations. This underscores the robustness of the TTC system in handling various interactions, where precise locations may not always be guaranteed.
Figure 8: Robustness over positional perturbations. We evaluate TTC-MonoDETR under controlled perturbations during test-time inference. Specifically, we introduce translation offsets to the ground-truth 2D centers with maximum ratios of 0%, 10%, 20%, 30%, and 40%.
Robustness over Visual Prompt Collection Frequency. In practical deployment, human feedback and open-vocabulary 2D detectors (e.g., Grounding DINO) are often delayed, making it infeasible to provide corrections for every missed detection in real time. To evaluate the robustness of TTC under varying feedback frequencies, we design experiments that simulate different feedback intervals.
TABLE 4: mAP comparisons under different feedback collection intervals. “N=0” means feedback is collected every frame, while “N=2” means every 2 frames.
Feedback Interval NN 0 2 4 6
TTC-BEVFormer 47.8 47.5 47.3 46.9
TTC-Sparse4Dv2 53.5 53.1 52.4 51.8
TTC-MV2D 51.0 50.7 50.2 49.6
As shown in Table˜4, we update the prompt buffer at different frame intervals to mimic varying levels of feedback availability. Despite reducing the update frequency (from every frame to every 6 frames), TTC maintains strong error correction performance, demonstrating its resilience to infrequent feedback. This suggests that TTC can effectively operate in real-world scenarios where continuous supervision is impractical.
Robustness over Noisy Visual Prompts. To evaluate robustness against completely misleading prompts, we augment the fixed prompt buffer used in Table˜3 with two animal exemplars (one dog, one kangaroo; prompt images in Appendix Figure˜21) in addition to vehicle prompts. These classes are virtually absent in nuScenes imagery and thus serve as off-topic (noisy) prompts. We evaluate the Car category using TTC-MonoDETR from Table˜3 (trained with 20% vehicle labels), comparing runs with and without the added animal prompts. As shown in Table˜5, adding these off-topic prompts do not degrade accuracy, further demonstrating TTC’s robustness to noisy prompts.
TABLE 5: Robustness to noisy visual prompts. Comparison of fixed-buffer TTC-MonoDETR with and without off-topic animal prompts (dog & kangaroo). Evaluation follow the protocol of Table˜3.
Prompt Type mAP $\uparrow$ EDS $\uparrow$
Vehicle 20.9 45.3
Vehicle + Animal 20.9 45.3
Robustness over Rapidly Changing Prompts. We test TTC under fast ego–motion and frequent occlusions, where prompts appear/disappear quickly. To emulate this regime, we uniformly sub–sample each video by 50%, effectively doubling apparent scene speed and shortening object dwell time, while keeping the 5-frame feedback interval and all thresholds/buffer rules identical to Table˜1. As summarized in Table˜6, the gains over the base detectors remain consistently positive (mAP/EDS and FN/frame), albeit smaller than in the full–frame setting. This indicates that TTC is robust to rapidly changing prompts.
TABLE 6: Robustness to rapidly changing prompts. We choose BEVFormer as the base detector.
Method mAP $\uparrow$ EDS $\uparrow$ FP/frame $\downarrow$ FN/frame $\downarrow$
BEVFormer 40.2 49.6 34.9 1.9
TTC-BEVFormer 45.9 (+5.7) 51.9 (+2.3) 35.0 1.2
4.4 Additional Comparisons and System Analysis
Beyond evaluating the effectiveness of TTC and the robustness of visual prompts, we conduct additional comparisons to clarify its advantages over alternative approaches: (i) comparison with an SOT-based online error–correction baseline and (ii) contrast with offline fine–tuning driven by test–time 2D feedback. We also provide system–level analyses: (iii) efficiency in terms of model size and inference latency; (iv) cross–dataset transfer ability; (v) feasibility of fully automated feedback based on 2D–3D misalignment; and (vi) FP trade-off induced by visual prompts.
Comparison with SOT-based Online Error Correction. An intuitive way to integrate online error correction is to combine a 2D tracking algorithm with a 3D detection head. To explore this possibility, we implement a baseline using ATOM [danelljan2019atom] for 2D tracking and MV2D [wang2023object] as the 3D object detector. MV2D leverages ATOM’s tracking results of missing objects from previous frames to initialize queries, mimicking an error correction mechanism. In this setup, we perform human intervention every 5 frames to provide a list of missing objects along with their 2D bounding boxes for ATOM as the reference. For a fair comparison, the visual prompts of TTC-MV2D are derived from user-drawn bounding boxes every 5 frames. As shown in Table˜7, TTC, equipped with the OA module, significantly outperforms this SOT-based approach in error correction. This highlights the limitations of simply combining SOT with a 3D detector and underscores the necessity of a dedicated online correction framework.
Method mAP (%) $\uparrow$ EDS (%) $\uparrow$
MV2D 42.6 51.9
SOT + MV2D 42.1 51.2
TTC-MV2D 50.0 57.3
Comparison with Offline Fine-tuning on Test-time Feedback. An alternative way to leverage auxiliary feedback is offline fine-tuning, where missing objects are collected during inference and later used to fine-tune the model with test-time 2D ground truth. While this enables adaptation, it introduces inherent delays compared to real-time correction. We evaluate this strategy by fine-tuning MV2D [wang2023object], which relies on 2D detection for 3D object inference and can incorporate 2D feedback. In this experiment, we collect human 2D feedback at different intervals and compare MV2D’s fine-tuned performance with TTC-MV2D, where prompts are updated at the same intervals. Results in Table˜8 show that TTC-MV2D consistently outperforms the fine-tuned MV2D with 2D weak feedback. This demonstrates its effectiveness of TTC in real-time streaming applications with only weak feedback feasible.
TABLE 8: mAP comparison between TTC-MV2D and MV2D fine-tuned with 2D feedback annotations. “N=0” means collecting feedback at every frame, while “N=2” means every 2 frames. TTC-MV2D consistently outperforms the offline fine-tuned MV2D.
Feedback Interval NN 0 2 4 6
MV2D + Offline fine-tune 43.4 42.7 41.0 39.9
TTC-MV2D 51.0 50.7 50.2 49.6
Efficiency Analysis: Model Size and Latency. For online deployment, a system must balance accuracy and efficiency. We analyze TTC ’s model size and inference latency, particularly in comparison with recent LLM-based promptable 3D detection methods [chen2024ll3da, huang2023embodied].
Method LL3DA MonoDETR TTC-MonoDETR
#Params $\downarrow$ 118M 68M 80M
FPS (Hz) $\uparrow$ 0.42 11.1 9.1
Latency in promptable systems arises from two sources: (i) response time in providing prompts and (ii) the system’s inference speed. While the first factor is unavoidable, the second is crucial for real-time applications. As shown in Table˜9, LLM-based methods like LL3DA [chen2024ll3da] exhibit significantly higher latency, making them impractical for autonomous driving. In contrast, TTC introduces minimal overhead compared to its base detector while maintaining real-time performance, confirming its suitability for online deployment in safety-critical scenarios.
Cross-dataset Transfer Analysis. We complement the results with a cross-dataset study from nuScenes to Waymo. Using MonoDETR as the base detector, Table˜10 shows that direct transfer causes a severe collapse (42.6 $\to$ 0.7 mAP). Accordingly, directly applying TTC brings only a marginal gain of +0.3 mAP. As discussed in Cross-Dataset Sensor Alignment [zheng2023cross], for vision-only 3D detectors the dominant factor is the mismatch in camera intrinsics/extrinsics: when moving from nuScenes to Waymo, objects at comparable 3D distance/shape occupy markedly different 2D pixel scales due to focal-length disparities and rig configurations (Appendix Figure˜11), which induces depth bias.
To decouple this confound, the intrinsic synchronization (K-sync) strategy [zheng2023cross] is adopted: images are resized and cropped to normalize focal lengths prior to inference. Under K-sync, TTC yields a substantial improvement over the K-sync–only baseline on Waymo ( +3.3 mAP ), indicating that once intrinsics are aligned, the proposed prediction-level rectification generalizes across datasets without updating model weights (Table˜10).
TABLE 10: mAP results of cross-dataset validation. The model is trained on nuScenes and evaluated on the nuScenes and Waymo. MonoDETR is used as the base detector.
Method nuScenes Waymo
MonoDETR 42.6 0.7
TTC-MonoDETR 45.0 1.0
K-sync + MonoDETR 42.8 30.8
K-sync + TTC-MonoDETR 45.8 (+3.0) 34.1 (+3.3)
Fully Automated Feedback Analysis. To address the practical difficulty of obtaining clicks/boxes while driving, we evaluate a most realistic deployment mode in which all prompts are generated automatically (no human inputs). We harvest feedback from the discrepancy between a 2D detector and the image reprojection of the current 3D predictions. Concretely, we consider two 2D detectors: (i) off-the-shelf Grounding DINO [liu2024grounding], and (ii) a YOLOv8 [ultralytics_yolov8] model retrained on the nuScenes 2D taxonomy. For each frame, every 2D box is compared against the projected 3D boxes; unmatched cases (IoU <0.5<0.5; optionally gated by reprojection residual) are converted into instance-specific visual prompts by cropping the corresponding image patch and enqueuing it into the buffer. Under the same in-domain protocol as Table˜1, automated prompts yield consistent improvements in mAP/EDS (Table˜11).
TABLE 11: Automated feedback results. 2D–3D misalignment is computed between 2D boxes (Grounding DINO / YOLOv8-nuScenes) and the image reprojection of 3D predictions; unmatched 2D boxes are converted into visual prompts.
Method mAP $\uparrow$ EDS $\uparrow$
BEVFormer 41.1 50.2
TTC-BEVFormer (GDINO) 41.8 (+0.7) 50.4 (+0.2)
TTC-BEVFormer (YOLOv8) 42.4 (+1.3) 50.8 (+0.6)
FP Impact Analysis. Because visual prompts introduce additional hypotheses at inference, a natural concern is whether they inflate false alarms. We therefore quantify the effect of TTC on false positives in Table˜1. Concretely, we report FP/frame and FN/frame over the full validation stream, where $\text{FP/frame}=\tfrac{\text{FP}_{\text{total}}}{\text{frames}}$ and $\text{FN/frame}=\tfrac{\text{FN}_{\text{total}}}{\text{frames}}$. A detection is matched using the same center–distance protocol as mAP calculation (averaged over $\{0.5,1.0,2.0,4.0\}$ m). Across diverse backbones, TTC markedly reduces FN/frame, while FP/frame changes only marginally ($\leq\!+0.2$), indicating that the recall gains do not come at the cost of inflating false alarms. We also evaluate a deployment-oriented setting with a confidence threshold of $0.3$; the FP/frame is not statistically significant before vs. after applying TTC (see Appendix Table˜15).
Why FP remains controlled. As stated in our method, each prompt spawns up to $N$ hypotheses (N=4N{=}4). For clarity, we define a valid prompt is one whose associated object is present in the current scene, whereas an invalid prompt is stale or contradictory (the object has left the scene or has no corresponding instance). On TTC–BEVFormer, we observe on average only $3.9$ visual prompts per frame; each valid prompt yields 1.6 above-threshold candidates on average, and each invalid prompt yields 0.3, indicating that visual prompts do not flood the detector and that invalid prompts are usually low-confidence. For above-threshold candidates, the outcomes are: (i) correct match to the intended object (desired); (ii) match to a visually similar object that has already been detected—suppressed by NMS, so no extra box appears; (iii) match to a similar object that is not detected—this becomes a TP for that object rather than an FP; (iv) fall on an empty region but overlap an FP from the dense query results—removed by NMS, thus not affecting the final FP count; (v) fall on an empty region with no overlap—this is a true FP, but it is rare (0.2 per frame in our statistics). In practice, (i) and (ii) dominate, while (v) is negligible; consequently, FP/frame stays near baseline as FN/frame drops (Appendix Figure˜12).
4.5 Ablation Studies
To better understand the key design choices in TTC, we conduct a series of ablation studies. Specifically, we investigate: (i) the contribution of core alignment components such as similarity and localization losses, (ii) the effect of addressing instance ambiguity through a one-to-$N$ mapping strategy, (iii) the capability of visual prompt alignment in achieving instance-level object differentiation, and (iv) the dynamic behavior of the visual prompt buffer during online inference. These studies provide deeper insights into the effectiveness and robustness of TTC in leveraging visual prompts for online 3D detection.
Effect of Components in Visual Prompt Alignment.
TABLE 12: Effect of Visual Prompt Alignment. We select TTC-MonoDETR for this experiment.
Sim. Loss Loc. Loss mAP (%) EDS (%)
- - 32.6 44.2
$\surd$ - 39.4 48.5
$\surd$ $\surd$ 43.3 52.3
Visual prompt alignment aims to localize objects via visual prompts in input images. We now evaluate its components with TTC-MonoDETR. Table˜12 presents experiments to verify the core components of this alignment, including similarity loss (Focal and Dice loss) and localization loss (supervising visual prompt localization, $\mathcal{X}_{v}$). While the alignment can be implicitly learned by attention mechanisms in the transformer decoder, incorporating explicit similarity supervision brings improvements of 6.8% mAP and 4.3% EDS. Further utilizing the position loss and one-to-$N$ mapping (to address instance ambiguity) boosts the performance to 43.3% mAP and 52.3% EDS. These results prove this alignment operation is a critical component enabling the TTC detectors to effectively detect target objects via visual prompts.
Figure 9: Visual prompt showcase for Table˜13.
Group mAP (%) $\uparrow$ EDS (%) $\uparrow$
a. 20.9 45.3
b. 13.8 42.2
Instance Ambiguity in Visual Prompt Alignment.
TABLE 14: Effect of the number of predicted positions $N$ of visual prompt alignment. We select TTC-MonoDETR for this experiment.
No. of Position Predictions mAP (%) $\uparrow$ Recall (%) $\uparrow$
1 39.9 62.1
4 43.3 69.1
8 43.0 69.4
To solve the instance ambiguity issue, we propose to regress $N$ positions of each visual prompt when performing the visual prompt alignment. This retrieves all objects with similar visual contents. In this study, we validate the effectiveness of this design by conducting ablation studies on the number of $N$. As listed in Table˜14, when the number of position predictions $N$ equals 1, which means a one-to-one mapping for each visual prompt, the mAP and recall rate are 39.9% and 62.1%, respectively. Then, if we increase the $N$ to 4, effectively a one-to-four mapping, we obtain an mAP of 43.3% and a recall rate of 69.1%. This represents a 7% improvement in the recall rate, demonstrating that instance ambiguity is an important challenge in visual prompt-based detection, and the proposed one-to-$N$ mapping solution effectively addresses this issue.
Effect of Visual Prompt Alignment on Instance Awareness. Despite demonstrating that visual prompt alignment can enhance system performance, we remain uncertain whether the alignment can distinguish different objects based on visual prompts for instance-level matching.
To investigate, we conduct an additional experiment, based on a similar setting with Table˜3 (TTC-MonoDETR trained on 20% vehicle annotations and freeze the prompt buffer with predefined web prompts when inference.), but fix the visual prompt buffer with images of black sedans solely (Figure˜9). As shown in Table˜13, as “Car” objects in nuScenes contain various types and colors, solely using black sedans as visual prompts leads to a 7.1% mAP drop. This underscores that the alignment operation effectively differentiates objects based on visual prompts, achieving instance-level matching and detection.
Statistics of Visual Prompt Buffer during Online Inference. We design the visual prompt buffer to store missed objects during inference with the video stream and introduce a “dequeue” mechanism to prevent the buffer from growing indefinitely. In this ablation study, we analyze the dynamic buffer size, as well as the number of enqueued and dequeued in each frame, to illustrate the behavior of the visual prompt buffer during the online operation of TTC. As shown in Figure˜10, the visual prompt buffer exhibits three distinct behaviors during online inference in each nuScenes video clip: increasing, steady, and decreasing. In initial frames, many traffic cones are queued into the buffer due to the poor performance of deployed offline 3D detectors on cone objects. The buffer size thus grows quickly in initial frames to store visual prompts of missed objects for online rectification (Frames #0 to #4). The buffer size then stabilizes as the online detector consistently detects and tracks all objects of interest (Frames #4 to #20). Further, as the ego vehicle drives out of the scene, many previously enqueued objects no longer exist and are thus removed from the buffer automatically (Frames #20 to #40). The buffer finally becomes empty as the scene changes. This demonstrates the effectiveness of the visual prompt buffer, which consistently stores missed objects during online inference and corrects online errors. This dynamic behavior, exhibiting increasing, steady, and decreasing phases, highlights itself to manage stored visual prompts for robust 3D object detection performance throughout the online inference with the balance between latency and accuracy. We also report several compact, deployment-oriented statistics over the full validation stream on TTC-BEVFormer: (i) the largest buffer size observed over the entire stream is only 65, with an average of 3.9; (ii) after an object leaves the scene, only 10.3% of its prompts remain in the buffer; (iii) only 17.2% of those left objects in (ii) produce erroneous final box after standard post-processing (confidence filtering + NMS). Together, these results indicate that bounded buffer growth, timely removal of outdated prompts, and negligible impact on final predictions—consistent with the dynamics in Figure˜10.
@@ 5 Conclusion
For model design, while allowing auxiliary information to provide reliability, it introduces potential risks. Requiring frequent feedback, especially human feedback, may be challenging in current L3 autonomous driving paradigms. Though reducing the frequency of feedback can alleviate this pressure without compromising performance, there is room for improvement. Future work could integrate robust tracking methods or leverage advanced hardware, such as eye-tracking, to simplify the feedback collection.
@@ Acknowledgement
This work was supported by National Key R&D Program of China (2022ZD0160104), NSFC (62206172), and Shanghai Committee of Science and Technology (23YF1462000). This work is in part supported by the JC STEM Lab of Autonomous Intelligent Systems funded by The Hong Kong Jockey Club Charities Trust. We thank team members from OpenDriveLab for valuable feedback along the project. Special thanks to Chengen Xie for providing data support, and Jiahui Fu for sharing insights on MV2D. We also sincerely thank Qingsong Yao for providing writing guidance and suggestions on experimental design.
@@ Appendix A Discussions
We emphasize, in this paper, we do not propose the online system to replace the traditional offline data loop. As illustrated in Figure˜2, these two systems address different aspects of autonomous driving. The offline system remains crucial for enhancing the capabilities of the base perception model through development; while our online TTC system further enables the deployed frozen model in vehicles to promptly rectify dangerous driving behaviors caused by unrecognized objects on the road. With improved offline-trained detectors, the TTC can effectively correct more online errors during test-time inference without re-training, as detailed in Table˜1.
@@
Compared to text prompts, visual prompts provide a more natural and accurate description of the target object. In contrast, verbal descriptions can be ambiguous to convey instance-level features, leading to an inaccurate understanding of the missing objects. Second, text promptable models are often combined with LLMs with high latency and are thus unavailable for autonomous driving deployment, as discussed in Section˜4.4.
Box and point prompts are less convenient than visual prompts when dealing with stream data. If missing occurs, these single-frame prompts require feedback at every frame, which is unfeasible in real-world applications. Compared to box and point prompts, visual prompts are robust across different scenes and timestamps, one single-frame visual prompt is enough for detecting and tracking in later frames. Furthermore, visual prompts enable 3D detection with pre-defined visual descriptions of target objects, regardless of the sources, styles, poses, etc, as discussed in Section˜4.3.
The introduction of novel visual prompts enables real-time, accurate, and continuous error correction of streaming inputs.
@@
Q4: What are potential applications and future directions of TTC?